OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of two x plus three, over five.

OpenStudy (anonymous):

f−1(x) = The quantity of five x plus three, over two f−1(x) = The quantity of five x minus three, over two f−1(x) = Five over the quantity of two x plus three f−1(x) = Five over the quantity of three minus two x